Problem 3-19 (Algo) Multiple Predetermined Overhead Rates; Applying Overhead (LO3-1, LO3-2, LO3-4] High Desert Potteryworks mRequired: 1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department. 2. ComputeComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req ЗА Req 3B Compute the total overhead costReq 1 Reg 2 Req ЗА Req 3B What would be the total manufacturing cost recorded for Job 205? (Round Predetermined overhead ratReg 1 Req 2 Req ЗА Req 3B If the job contained 24 units, what would be the unit product cost? (Round Predetermined overhead
Answer #1

1) Molding department = (178200/81000+3) = 5.20 per MH

Painting department = (476080/54100+5) = 13.80 per DLH

2) Total overhead cost = (380*5.20+127*13.80) = 3728.60 or 3729

3a) Total manufacturing cost

Molding painting Total
Direct material 950 1300 2250
Direct labor 700 990 1690
Overhead 1976 1753 3729
Total 3626 4043 7669

3b) Unit product cost = 7669/24 = 319.54